OpenStudy (chaotic_butterflies):

3000 = 2000 (2.035)^2t How do I solve for t?

OpenStudy (aveline):

Start out by dividing each side by 2000

OpenStudy (chaotic_butterflies):

3/2 = (2.035)^2t

OpenStudy (aveline):

log(3/2)/log(2.035)=2t

OpenStudy (chaotic_butterflies):

Gotcha, thanks!